# Evidence for itinerant electron-local moment interaction in Li-doped -MnTe Tingjun Zhang

### v1

@@ Abstract
We use inelastic neutron scattering to study the impact of Li doping on the semiconducting altermagnet -MnTe. Introducing Li results in a spin reorientation from in-plane to out-of-plane and increases the density of itinerant carriers. While the spin waves in Li-doped -MnTe remain largely Heisenberg-like, there is significant spin wave broadening across the entire Brillouin zone, signaling enhanced magnon damping induced by itinerant carriers. By extracting the local dynamic susceptibility and applying the total moment sum rule, we find that both undoped and Li-doped -MnTe
exhibit the full expected Mn2+ local moment of
with .
These results demonstrate that Li-doped -MnTe hosts robust local-moment magnetism whose interactions are mediated via Ruderman-Kittel-Kasuya-Yosida-type interactions, highlighting the importance of itinerant carriers in magneto-transport and spin dynamic properties of altermagnets.
@@
the collective motion of localized moments in the magnetic ordered state is spin waves (magnons) characterized by linear spin wave theories (LSWT) that ignore all terms of order higher than quadratic [WOS:000207403700002, PhysRev.58.1098, PhysRev.117.117, 10.1093/oso/9780198862314.001.0001]. Since magnon-magnon interactions are negligible [RevModPhys.85.219], the lifetime of magnons as measured by the energy width of spin wave spectra should be infinite and only limited by
@@
the instrumental resolution, as seen in ferromagnetic EuO [PhysRevB.14.4923] and collinear antiferromagnetic Rb2MnF4 [PhysRevLett.111.017204]. The dispersion of spin waves determines the nearest-neighbor (NN), next-nearest-neighbor (NNN), etc. magnetic exchange couplings, and, therefore, the magnetic ordering temperature, regardless of the ordered moment direction controlled by the spin-orbit coupling (SOC)
@@
[10.1093/oso/9780198862314.001.0001]. For classical spin systems (i.e.,) without strong quantum fluctuations, geometric frustration, or appreciable spin-lattice coupling,
magnon decay processes are generally not expected in insulating magnets [RevModPhys.85.219].
In metallic magnets, spin wave broadening can arise from a variety of interactions with conduction electrons. In itinerant systems, magnons can enter the Stoner particle–hole continuum at high energies, where they experience strong Landau damping [Stoner1936, Stoner1938, PhysRev.52.198, RevModPhys.25.211]. In local-moment metals, low-energy spin-flip scattering of conduction electrons can produce Korringa-type damping in the long-wavelength limit.
The phenomenological Landau–Lifshitz–Gilbert (LLG) model [Landau1935, Gilbert1955] captures these long-wavelength relaxation processes, but these typically yield an extremely small intrinsic contribution to damping in three-dimensional bulk magnets. Even in the canonical metallic ferromagnet -Fe, spin waves remain resolution-limited up to energies of order 200 meV, only broadening upon entering the Stoner continuum [PhysRevB.11.2624, 10.1063/1.348814, brookes_spin_2020]. Thus, significant spin wave damping is often weak or difficult to observe, and in systems with mixed local moment / itinerant character, its magnitude can depend sensitively on the structure of the Fermi surface.
In the semiconducting altermagnet -MnTe, the opposite magnetic sublattices of Mn2+ are connected by symmetry operations that are not inversion () or translation () symmetric due to the specific local crystal environment [Fig. 1(a)]. This leads to alternating spin splitting of electronic energy bands, arising from exchange coupling rather than SOC [smejkal_CrystalTimereversalSymmetry_2020, smejkal_ConventionalFerromagnetismAntiferromagnetism_2022, smejkal_EmergingResearchLandscape_2022]. This is in sharp contrast to most collinear antiferromagnets, where the combination of time-reversal symmetry () and or symmetries between opposite magnetic sublattices enforces the degeneracy of electronic bands throughout the entire Brillouin zone (BZ) in the limit of zero SOC [10.1093/oso/9780198862314.001.0001]. Indeed, several phenomena commonly associated with the breaking of -symmetry – spin-split electronic bands [din_unconventional_2025], anomalous Hall effects (AHE) [gonzalez_betancourt_spontaneous_2023, kluczyk_coexistence_2024, PhysRevLett.134.086701, liu_StraintunableAnomalousHall_2025, smolenski2025straintunabilitymultipolarberrycurvature],
and magnetic circular dichroism [hariki_x-ray_2024] – have been experimentally realized in altermagnetic materials [barman_2021_2021, flebus_2024_2024].
In addition to momentum-dependent electronic energy band splitting, altermagnetic symmetry also gives rise to a momentum-dependent energy splitting of chiral magnons which is non-relativistic in origin and holds even in the limit of symmetric exchange interactions [liu_chiral_2024].
Figure 1:
@@
(a,b) Parallel and top views of the lattice and magnetic structures of -MnTe, highlighting magnetic moment directions and exchange couplings. The dashed lines indicate the altermagnetic exchange interactions and .
(c,e) Schematics of magnetic domain configurations in undoped and Li-doped -MnTe, respectively. The edges of the hexagons are parallel to and equivalent crystallographic directions in panel (b).
(d) Parallel view of the lattice and magnetic structures of Li-doped -MnTe, illustrating Li atoms (occupancy exaggerated) occupying Te-site defects.
(f) Magnetization of Li-doped -MnTe versus temperature .
(g) Longitudinal resistivity versus temperature for undoped and Li-doped -MnTe.
(h) Transverse resistivity of Li-doped MnTe.
(i) Carrier density and mobility of undoped and Li-doped -MnTe.
@@
(j,k) Calculated spin waves of undoped -MnTe for three domains within (j) and outside (k) the nodal planes.
The calculations use exchange parameters from Ref. [liu_chiral_2024].
@@
The inset in (k) displays the intensity contributions from the three domains at the point of altermagnetic splitting. The solid and dashed lines denote different branches of the spin waves, and the vertical dashed line separates the splitting dominated by single-ion anisotropy () and altermagnetism ().
@@
(l,m) Calculated spin waves of Li-doped -MnTe, with the single-ion anisotropy replaced by the value reported in Ref. [yumnam_MagnonGapTuning_2024].
@@
An attractive aspect of -MnTe in the context of spintronics is the tunability of its altermagnetic ground state, which can be modified by
chemical substitution [moseley_giant_2022, yumnam_MagnonGapTuning_2024],
uniaxial strain [PhysRevLett.134.086701, liu_StraintunableAnomalousHall_2025, smolenski2025straintunabilitymultipolarberrycurvature],
@@
and external applied fields [kriegner_multiple-stable_2016, kriegner_magnetic_2017, amin_nanoscale_2024, gonzalez_betancourt_anisotropic_2024, dzian_antiferromagnetic_2025]. -MnTe adopts a layered -type magnetic order below K [kunitomiNeutronDiffractionStudy1964, efremdsaLowtemperatureNeutronDiffraction2005, szuszkiewiczNeutronScatteringStudy2005, kriegnerMagneticAnisotropyAntiferromagnetic2017, szuszkiewiczSpinwaveMeasurementsHexagonal2006], where Mn2+ moments are oriented along the NNN Mn-Mn direction [Fig. 1(a,b)] [liu_StraintunableAnomalousHall_2025].
@@
The anisotropy of the Mn2+ moments can be strongly modified from in-plane to out-of-plane by less than 1% Li doping in -MnTe, which also introduces charge carriers into the system [Fig. 1(d)] [yumnam_MagnonGapTuning_2024].
This spin reorientation removes the three-fold degeneracy in the magnetic channel, and effectively provides another pathway to achieving a single domain response without modifying the symmetries required for altermagnetism [Fig. 1(e)] [smejkal_CrystalTimereversalSymmetry_2020, smejkal_ConventionalFerromagnetismAntiferromagnetism_2022, smejkal_EmergingResearchLandscape_2022].
Figure 2:
(a,b) INS spectra acquired at base temperature ( K) with an incident energy of meV for Li-doped [high-flux (HF) mode] and undoped -MnTe, shown along high-symmetry directions within the nodal plane.
@@
(c-e) Energy-axis line cuts at selected points marked in panel (a), including both HF and high-resolution (HR) modes for Li-doped -MnTe.
Dashed lines indicate Gaussian fits. Horizontal bars mark the instrumental energy resolution at the Gaussian peak position for each configuration. (f) Schematic of the Brillouin zone indicating high-symmetry points. The nodal planes are shown in gray, and the -A–L–M plane is highlighted in red.Figure 3:
(a,b) INS spectra along and (c,d) along the direction for undoped and Li-doped -MnTe.
(e) Schematic of the -A-L-M plane of the Brillouin zone, where splitting is expected to occur in undoped -MnTe.
@@
For Li-doped -MnTe, the nonzero values of and were fixed to the refined values from undoped -MnTe.
Data for Li-doped -MnTe was shifted vertically by 0.075 units for clarity.
(h,i) Wavevector dependence of the energy full-width at half-maxima (FWHM) near the zone center along -A and -M directions, respectively.
Understanding the impact of charge carriers on this prototypical altermagnetic ground state and its excitations is important.
The AHE, for example, has been shown to be highly sample-dependent based on the magnitude of the experimental charge gap, revealing a narrow window of in-gap impurity states for which the AHE is observable in -MnTe [liu_StraintunableAnomalousHall_2025].
The chiral splitting of the magnons is driven by the difference in the exchange pathways and which, while equal in distance, are differentiated by the specific pathway through Te orbitals [Fig. 1(a)] [liu_chiral_2024].
For spin waves in Li-doped -MnTe, while the spin reorientation is expected to renormalize the domain contributions and the spin gap induced by SOC [Fig. 1(j-m)], it remains unclear how the addition of charge carriers and the associated structural substitution of Li into the lattice impact the spin dynamics which underlie the functionality of the altermagnetism in this system.
To address this, we investigate the effects of 3% Li doping on the spin waves of -MnTe via single-crystal inelastic neutron scattering (INS) measurements, and present a quantitative comparison of these results with those from undoped -MnTe.
While spin waves in both undoped and Li-doped -MnTe can be well-described within a local-moment framework, the addition of charge carriers via Li doping modifies the spin anisotropy gap through SOC [yumnam_MagnonGapTuning_2024], induces magnon damping at all energies, and obscures the chiral magnon splitting seen in undoped -MnTe – all via interactions with itinerant electrons.
Therefore, itinerant electron-local moment interactions in
-MnTe are essential not only in determining the AHE
[liu_StraintunableAnomalousHall_2025], but also for understanding spin dynamics in a system dominated by Mn2+ local moments.
To determine the position and occupancy of Li atoms, we performed single crystal neutron diffraction measurements and refined several structural models against the experimental data. The results indicate an approximately 4% Te-site deficiency in both samples (Tables S5–S10).
In the Li-doped sample, allowing Li to occupy the Te site yields a stable refined Li occupancy of 3% at both 90 K and 340 K, whereas placing Li on the site or on tested interstitial positions drives the Li occupancy to values that are effectively zero and does not improve the agreement factors (Table S3).
The temperature dependence of the magnetization of the 3% Li-doped sample is consistent with a Néel vector along the axis [Fig. 1(f)] [yumnam_MagnonGapTuning_2024], while the resistivity drops by two orders of magnitude relative to undoped -MnTe [Fig. 1(g)].
Notably, remains unchanged at this doping level, reminiscent of the robust antiferromagnetic ground state observed in K-doped BaMn2As2 [lamsal_persistence_2013]. This suggests that the effective exchange energy scale between Mn2+ local moments remains similarly unchanged.
The extracted carrier density and positive Hall coefficient indicates that holes are the dominant carrier in both systems, and Li doping enhances their contribution to conduction [Fig. 1(h,i)].
To understand the impact of Li-doping on the spin waves of -MnTe,
we conducted INS measurements at the SEQUOIA spectrometer at the Spallation Neutron Source for undoped -MnTe and at the MAPS Spectrometer at the ISIS Neutron Source for Li-doped -MnTe.
Importantly, both experiments were carried out on a single piece of single crystal with mosaic less than 1 degree using a nearly identical resolution and normalized to absolute units in order to ensure a quantitative comparison of the experimental spectra [10.1063/1.4818323].
The chiral magnon splitting for pure -MnTe is observed in the -A-L-M plane of BZ, highlighted in red in Fig. 2f, while spin waves remain degenerate in the nodal planes shaded in gray [liu_chiral_2024].
We first examine spin waves along directions in the nodal planes [Fig. 2 (a,b)], which reveal three differences induced by Li doping:
(3) whereas spin waves in -MnTe are resolution-limited in energy, the linewidth in the Li-doped case is substantially broadened [Fig. 2(c-e), S2].
Whereas (2) can be understood as the addition of indirect exchange via Ruderman-Kittel-Kasuya-Yosida (RKKY)-type interactions mediated by itinerant carriers, the -independent damping implied by (3) may have several possible physical origins for local moment metals.
First, the non-relativistic spin splitting in the electronic bands in altermagnets may allow intra-band spin-flip scattering along certain directions in reciprocal space, giving rise to strongly momentum-dependent magnon broadening.
However, spin waves in Li-doped -MnTe show a relatively -independent broadening across the BZ (including nodal planes) and at all energies (Figs. 3(h,i), S2). This stands in sharp contrast to itinerant magnets with comparable metallicity such as MnSi [petrova_behavior_2006, ishikawa_magnetic_1977, chen_unconventional_2020], FeGe [teng_spin-charge-lattice_2024], and -Fe [PhysRevB.11.2624, 10.1063/1.348814, brookes_spin_2020], where resolution-limited spin waves persist up to the appearance of a Stoner continuum.
Second, the quasi-random substitution of Li+ within the lattice can generate a distribution of local exchange coupling strengths mediated by RKKY interactions, and this exchange disorder can manifest as a smeared energy linewidth reflecting a superposition of the local exchange couplings.
We note, however, that the level of disorder between our samples of undoped and Li-doped -MnTe are nearly comparable.
Further, at a doping level of 3% Li per formula unit, there is only a 15% probability that a given Mn2+ would have Li+ as a NN [Fig. 1(b)], suggesting that a purely disorder-driven effect is insufficient to describe the strong renormalization of the spin wave lifetime.
In electron-doped iron pnictides BaFe2-xNixAs2, doping levels up to do not modify the high-energy spin excitations [RevModPhys.87.855], much different from the spin wave broadening seen in the entire BZ of Li-doped -MnTe [Figs. 3(h,i), S2].
Lastly, magnons may decay into low-energy particle-hole excitations near the Fermi level, as observed in the Mn-based metallic magnets YbMnBi2 [sapkota_signatures_2020] and YbMnSb2 [hu_coupling_2023] where a low density of states at the Fermi level results in a -independent damping at all energies.
As a measure of the modifications to the spin Hamiltonian within a local moment framework, we fit the spin wave spectra of both compounds using LSWT based on a Heisenberg Hamiltonian with a single-ion anisotropy term (Eqn. S1). A positive (negative) corresponds to the easy-plane (easy-axis) anisotropy in the undoped (doped) system. Compared to undoped -MnTe, the NN exchange coupling in the Li-doped sample is reduced by 8%, consistent with the spin wave bandwidth reduction.
To quantify the effects of an RKKY-mediated exchange-disorder scenario, we assume the spectral weight as a Gaussian distribution with a finite half-width at half-maximum (HWHM) centered around the spin wave dispersion .
While the large linewidths near the point is primarily attributed to overlap between (non-)degenerate optical and acoustic modes, the linewidths at the BZ boundaries more accurately capture the underlying broadening (Fig. S2).
Accordingly, we employed three sets of fitted energy values (HWHM, , and HWHM) along the M–L and L–A directions, and used alone along other high-symmetry directions as input for LSWT calculations. These are denoted as “upper”, “center”, and “lower”, respectively, in Table S1. Among the exchange parameters, the further-neighbor interactions and the anisotropy term were found to be relatively insensitive to this spectral weight distribution. In contrast, the term exhibited a distribution range of meV, corresponding to a HWHM of about 1.5 meV in the spectral weight at the BZ boundary.
@@
We now inspect the spin wave spectra on the -A-L-M plane of the BZ, where splitting is expected for undoped -MnTe [Fig. 3(e)] [liu_chiral_2024]. Figure 3(a-d) show a comparison of spin wave spectra along the and directions for both systems.
@@
While splitting is clearly observed in undoped -MnTe [liu_chiral_2024], the broadened linewidth in Li-doped -MnTe makes the splitting more challenging to resolve.
@@
This is also reflected in constant energy cuts [Fig. 4(a-c)], where inspection of the -dependence of the ring-like spectral weight reveals an alternating bright/dim pattern around the ring, which is a direct result of the splitting in undoped -MnTe. We note that in Li-doped -MnTe a weak alternation is evident at 34.5 meV, which may signal a weak splitting which survives Li-doping [Fig. 4(c)].
@@
To more closely investigate the presence of splitting in the two systems, we performed constant- cuts near the expected splitting position in - space [Fig. 3(e-g)].
@@
While two peaks are resolvable in the line cuts for undoped -MnTe, the spectra for Li-doped -MnTe are evidently broader and only a single, broad peak is resolvable for most values.
Figure 4:
@@
(a-c) Constant-energy maps of the INS spectra in the plane for undoped (left) and Li-doped -MnTe (right) between 33.5 meV and 34.5 meV. The BZ is outlined by a thin grey line. The observed splitting in undoped -MnTe is indicated by red arrows. (d) Energy dependence of the local dynamic susceptibility extracted from INS data, with effective moments obtained from the total moment sum rule.
@@
To understand whether the observed dispersion and lineshape for Li-doped -MnTe can be explained by a persistent splitting which is simply broadened by the addition of itinerant carriers, we fit the data to two distinct LSWT models with refined , , : (1) with and zero, where splitting is necessarily absent; (2) with non-zero difference fixed at the value refined for undoped -MnTe. and were allowed to refine under this constraint [Fig. 3(f,g)].
In the former scenario, which neglects the splitting allowed by the altermagnetic symmetry, the energy of the spin waves in Li-doped -MnTe are underestimated near the expected splitting, and the experimental linewidth is substantially broader than predicted by the model.
@@
For the model with , we find improved agreement with experimental energies, further suggesting that the splitting in -MnTe may survive Li-doping.

### v2

@@ Abstract
We use inelastic neutron scattering (INS) and angle-resolved photoemission spectroscopy (ARPES) to study the impact of Li doping on the semiconducting altermagnet -MnTe.
Introducing Li results in a spin reorientation from in-plane to out-of-plane
direction and increases the density of itinerant carriers.
While our ARPES measurements do not indicate any notable doping-induced changes in the electronic band structure or the magnitude of the altermagnetic band splitting, our INS measurements reveal an abrupt carrier-induced decrease in the spin wave lifetime near the zone boundary at high energies. This finding is consistent with a new magnon decay channel driven by doping-induced subtle changes in the band structure and enhanced interactions between Mn2+ local moments and itinerant electrons. By extracting the local dynamic susceptibility from INS spectra and applying the total moment sum rule, we find that both undoped and Li-doped -MnTe exhibit the full expected Mn2+ local moment of with . These findings suggest that -MnTe hosts robust local-moment altermagnetism which shows a breakdown at high energies upon addition of itinerant carriers, highlighting the importance of carrier-spin coupling in magneto-transport and spin dynamic properties of altermagnets even in the dilute-carrier limit.
@@
the collective motion of localized moments in the magnetic ordered state is spin waves (magnons) characterized by linear spin wave theories (LSWT) that ignore all terms of order higher than quadratic [32, 33, 50, 7].
@@
Their dispersion determines the nearest-neighbor (NN), next-nearest-neighbor (NNN), etc. exchange couplings, and, therefore, the magnetic ordering temperature – regardless of the ordered moment direction controlled by the spin-orbit coupling (SOC)
[7].
Since magnon-magnon interactions are negligible [74], the magnon lifetime as measured by the energy width of spin wave spectra should be infinite and only limited by the instrumental resolution, as seen in ferromagnetic EuO [17] and collinear antiferromagnetic Rb2MnF4 [5].
@@
For classical spin systems (i.e.,) without strong quantum fluctuations, geometric frustration, or SOC,
magnon decay processes are generally not expected in insulating magnets [74].
@@
Figure 1: (a,b) Parallel view of the lattice and magnetic structures of MnTe and Li-MnTe, respectively, highlighting magnetic moment directions and exchange couplings. The dashed lines indicate the altermagnetic exchange interactions and . The Li occupancy at Te-site defects is exaggerated.
(c,d) Schematics of magnetic domain configurations MnTe and Li-MnTe, respectively.
(e) Magnetization of Li-MnTe versus temperature .
(f) Longitudinal resistivity versus for MnTe and Li-MnTe.
(g) Transverse resistivity of Li-MnTe.
(h) Carrier density and mobility of MnTe and Li-MnTe.
@@
(i,j) Calculated spin waves of MnTe for three domains within (j) and outside (k) the nodal planes.
The calculations use exchange parameters from Ref. [45].
@@
The inset in (k) displays the intensity contributions from the three domains at the position of altermagnetic splitting. The solid and dashed lines denote different branches of the spin waves, and the vertical dashed line separates the splitting dominated by single-ion anisotropy () and altermagnetism ().
@@
(k,l) Calculated spin waves of Li-MnTe, with the single-ion anisotropy replaced by the value reported in Ref. [73].
In metallic magnets, spin-wave excitations can decay through their coupling to conduction electrons. In strongly itinerant systems, magnons may enter the Stoner particle-hole continuum at elevated energies, where their lifetime is rapidly reduced by Landau damping arising from decay into electron-hole spin-flip excitations [65, 66, 60, 71]. In local-moment metals, the magnetic dispersion is primarily governed by the dipolar exchange interactions between well-defined local moments, while low-energy scattering of conduction electrons can produce Korringa-type damping that is most clearly manifested at long-wavelengths. The hydrodynamic form of such relaxation processes is commonly described phenomenologically within the Landau-Lifshitz-Gilbert framework [43, 26], which captures viscous damping of collective spin precession but does not by itself specify the underlying microscopic mechanism. In many three-dimensional metallic ferromagnets, this intrinsic damping remains relatively weak over a broad energy range; for example, in elemental -Fe spin waves are observed to remain essentially resolution-limited up to energies approaching the onset of the Stoner continuum near 200 meV [46, 52, 9]. Consequently, measurable spin-wave broadening in metallic magnets provides a sensitive probe of the degree of itinerancy and of the details of carrier-spin coupling encoded in the electronic structure near the Fermi level.
In the semiconducting altermagnet -MnTe (hereafter, MnTe), the opposite magnetic sublattices of Mn2+ are connected by symmetry operations that are not inversion () or translation () symmetric due to the specific local crystal environment [Fig. 1(a)]. This leads to alternating spin splitting of electronic energy bands, arising from exchange coupling rather than SOC [61, 62, 63]. This is in sharp contrast to most collinear antiferromagnets, where the combination of time-reversal symmetry () and or symmetries between opposite magnetic sublattices enforces the degeneracy of electronic bands throughout the entire Brillouin zone (BZ) in the limit of zero SOC [7]. Indeed, several phenomena commonly associated with the breaking of -symmetry – spin-split electronic bands [18], anomalous Hall effects (AHE) [27, 36, 6, 44, 64],
and magnetic circular dichroism [31] – have been experimentally realized in altermagnetic materials [4, 25].
In addition to momentum-dependent electronic energy band splitting, altermagnetic symmetry also gives rise to a momentum-dependent energy splitting of chiral magnons which is non-relativistic in origin and holds even in the limit of symmetric exchange interactions [45, 59, 67, 57, 24].
@@
An attractive aspect of MnTe in the context of spintronics is the tunability of its altermagnetic ground state, which can be modified by
chemical substitution [48, 73],
uniaxial strain [6, 44, 64],
@@
and external applied fields [40, 38, 2, 28, 19]. MnTe adopts a layered -type magnetic order below K [41, 20, 69, 39, 68], where Mn2+ moments are oriented in-plane along the NNN Mn-Mn direction [Fig. 1(a,c)] [44].
@@
The anisotropy of these Mn2+ moments can be strongly modified from in-plane to out-of-plane by less than 1% Li doping, which also introduces charge carriers into the system [Fig. 1(b,d)] [73].
This spin reorientation removes the three-fold degeneracy in the magnetic channel, and effectively provides another pathway to achieving a single domain response without modifying the symmetries required for altermagnetism [Fig. 1(d)] [61, 62, 63].
Figure 2: (a,b) INS spectra acquired at base temperature ( K) in HF mode with an incident energy of meV for Li-MnTe and MnTe, shown along high-symmetry directions within the nodal plane.
@@
(c) FWHM values obtained from Gaussian fits of the INS spectra in the nodal planes of the BZ for pure and Li-MnTe. Solid and dashed curves represent the energy-dependent instrumental resolutions from Fig. S1.
(d) FWHM values obtained by subtracting the FWHM of pristine MnTe from that of Li-MnTe. Vertical error bars represent 1 standard deviations and horizontal error bars indicate the integration range along the corresponding path.
(e) Energy-dependent line cuts at selected points along the -M path. The fitted FWHM values are listed within each sub-panel.
(f) Schematic of the BZ, where nodal planes are shaded in gray and the -A–L–M plane is shaded in red.
Filled symbols use Li-MnTe HF data and open symbols use Li-MnTe HR data in each panel.
Understanding the influence of charge carriers on the prototypical altermagnetic ground state of MnTe and on its collective excitations is important. The AHE, for example, has been shown to depend sensitively on sample-dependent variations in the charge gap, indicating that a narrow window of in-gap impurity states is required for its observation in MnTe [44]. At the microscopic level, the chiral splitting of the magnon branches originates from the inequivalence of the exchange pathways and which, despite having equal bond lengths, are distinguished by their distinct superexchange routes through Te orbitals [Fig. 1(a)] [45]. In Li-doped MnTe (hereafter, Li-MnTe), the extent to which the introduction of charge carriers and the associated substitutional disorder affect the spin dynamics that underpin altermagnetic functionality remains unresolved. To address this, we investigate the effects of Li doping on the magnetic excitations and electronic structure of Li-MnTe using single-crystal inelastic neutron scattering (INS) and angle-resolved photoemission spectroscopy (ARPES), respectively. We find that Li doping produces several notable changes, particularly a pronounced magnon damping near the zone boundary above meV, pointing to an increased role of interactions between itinerant electrons and Mn2+ local moments which are essential for understanding the emergence of the AHE [44] as well as the spin dynamics in this system.
First, we attempt to determine the position and occupancy of Li by refining several structural models against single crystal neutron diffraction data.
Our results indicate an approximately 4% Te-site deficiency in both MnTe and Li-MnTe (Tables S5–S11).
In Li-MnTe, allowing Li to occupy the Te site yields a stable refined Li occupancy of 3% at both 90 K and 340 K, whereas placing Li on the site or on tested interstitial positions drives the Li occupancy to values that are effectively zero and does not improve the agreement factors (Table S3).
The magnetization of the 3% Li-doped sample is consistent with a Néel vector along the axis [Fig. 1(e)] [73], while the resistivity drops by two orders of magnitude relative to MnTe [Fig. 1(f)].
Notably, remains unchanged at this doping level, reminiscent of the robust antiferromagnetic ground state observed in K-doped BaMn2As2 [42].
This suggests that the effective exchange energy scale between Mn2+ local moments remains similarly unchanged.
The extracted carrier density and positive Hall coefficient indicate that holes are the dominant carrier in both systems, and Li doping enhances their contribution to conduction [Fig. 1(g,h)].
To understand the impact of Li-doping on the spin waves of Li-MnTe,
we conducted time-of-flight INS measurements at the SEQUOIA spectrometer at the Spallation Neutron Source for pristine MnTe and at the MAPS Spectrometer at the ISIS Neutron Source for Li-MnTe. Measurements were performed on a single piece of single crystal with mosaic less than 1 degree, using nearly identical instrumental resolutions (Fig. S1) and normalized to absolute units in order to ensure a quantitative comparison of the experimental spectra [72].
In pristine MnTe, chiral magnon splitting appears in the -A-L-M plane, highlighted in red in Fig. 2f, while spin waves remain degenerate in the nodal planes shaded in gray [45].
We first examine spin waves within the nodal planes [Fig. 2 (a,b)], which reveal three differences induced by Li doping:
(3) whereas spin waves in pristine MnTe are resolution-limited throughout the BZ, the linewidth in Li-MnTe is notably larger near the zone boundary [Fig. 2(c-e)].
Whereas (2) can be understood as the addition of indirect exchange via Ruderman-Kittel-Kasuya-Yosida (RKKY)-type interactions mediated by itinerant carriers, the appearance of magnon damping implied by (3) may have several possible physical origins in metallic magnets, as previously discussed.
To distinguish between these scenarios, we performed energy-dependent Gaussian profile fits at constant to compare the energy full width at half-maximum (FWHM) between MnTe and Li-MnTe in high-flux (HF) mode, where the instrumental resolutions are essentially identical (Fig. S1). We also carried out high-resolution (HR) measurements for Li-MnTe, in order to differentiate between intrinsic and extrinsic broadening effects (Figs. S2, S3). While the linewidth at the point is dominated by the presence of a nearby optical mode [Fig. 1(i,k)], by moving slightly away from the point one can draw a quantitative comparison between the FWHM of Li-MnTe and MnTe. Taking the MnTe FWHM to represent the intrinsic linewidth, the difference profile (FWHM) between the two HF datasets demonstrates that the linewidth is nearly identical between the two samples away from the BZ boundary [Fig. 2(d)], with the exception of an artifact at intermediate (e.g., halfway along -M) which manifests as a hump in the HF dataset but is less pronounced in the HR dataset. In contrast, near the zone boundary we find a notable increase in the FWHM of Li-MnTe in both HF and HR modes, strongly suggesting that this is an intrinsic effect induced by Li doping. Examining the energy-dependence of this broadening reveals that its onset is limited to energies above meV, with slightly different onset energies at different (Fig. S4).
Figure 3: (a,b) INS spectra along and (c,d) along the direction for MnTe and Li-MnTe.
(e) Schematic of the -A-L-M plane of the BZ, where splitting is expected to occur in MnTe.
@@
For Li-MnTe, the nonzero values of and were fixed to the refined values from MnTe.
Data for Li-MnTe was shifted vertically by 0.075 units for clarity.
Such an abrupt reduction in the spin-wave lifetime is inconsistent with the smooth energy dependence expected for disorder-dominated damping [35, 8, 12]. In the dilute-defect limit, the magnon self-energy is governed by the single-impurity -matrix and scales linearly with defect concentration [35, 8]. For strong point-like scatterers, the resulting linewidth is expected to follow , where is the magnon energy, implying
that the linewidth increase should be proportional to
impurity level ; and should
gradually increase with rather than a sharp onset [12]. Comparable behavior is obtained in supercell and coherent-potential simulations of disordered Heisenberg magnets [10]. The abrupt broadening also differs qualitatively from the damping attributed to dynamical spin-lattice coupling in ferromagnet CrGeTe3 [11], which produces a more gradual energy evolution. Since spin-lattice coupling is not expected to change substantially upon light Li substitution, the pronounced high-energy damping in Li-doped MnTe instead points to an additional decay channel associated with increased itinerant character and interactions between itinerant electrons and local moments. A related finite linewidth broadening has been reported in low-carrier-density Mn-based metallic magnets such as YbMnBi2 [55] and YbMnSb2 [34].
As a measure of the modifications to the spin Hamiltonian within a local-moment framework, we fit the spin wave spectra of both compounds using LSWT based on a Heisenberg Hamiltonian with a single-ion anisotropy term (Eqn. S1). A positive (negative) corresponds to the easy-plane (easy-axis) anisotropy in the undoped (doped) system. Compared to MnTe, the NN exchange coupling in Li-MnTe is reduced by 8%, consistent with the spin wave bandwidth reduction. To quantify the possible impurity-induced exchange-disorder mechanism, we modeled the measured energy dependence of the spectral weight at each momentum transfer [] as a Gaussian profile with a finite HWHMFWHM/2. We then constructed three independent LSWT models for Li-MnTe based on distinct input dispersions along the BZ-boundary segments M-L and L-A, where the linewidth broadening is most pronounced. The reference (“center”) model employed the dispersion defined by the peak energies obtained directly from Gaussian fits to the experimental spectra. Two additional models were generated by systematically shifting these peak energies at each along these boundary paths by the corresponding linewidth, yielding upper and lower bounding dispersions defined by and , respectively. These three cases are denoted as “upper”, “center”, and “lower” in Table S1.
Among the exchange parameters, the further-neighbor interactions and the anisotropy term were found to be relatively insensitive to this spectral weight distribution. In contrast, the term exhibited a distribution range of meV, corresponding to a HWHM of about 1.5 meV in the spectral weight at the BZ boundary or about 10% of . Therefore, in Li-MnTe is 3 times larger than , again inconsistent with impurity-induced effect [35, 8, 12].
@@
We now inspect the spin wave spectra on the -A-L-M plane of the BZ, where splitting is expected in MnTe [Fig. 3(e)] [45]. Figure 3(a-d) show a comparison of spin wave spectra along the and directions for both systems.
@@
While splitting is clearly observed in MnTe [45], the broadened linewidth in Li-MnTe makes the splitting more challenging to resolve.
@@
This is also reflected in constant energy cuts [Fig. 4(a-c)], where inspection of the -dependence of the ring-like spectral weight reveals an alternating bright/dim pattern around the ring, which is a direct result of the splitting in MnTe. We note that in Li-MnTe a weak alternation is evident at 34.5 meV, which is consistent with a persistent splitting [Fig. 4(c)].
@@
To more closely investigate the presence of magnon splitting in the two systems, we performed constant- cuts near the expected splitting position in - space [Fig. 3(e-g)].
@@
While two peaks are resolvable in the line cuts for MnTe, the spectra for Li-MnTe are evidently broader and only a single, broad feature is resolvable for most values.
@@
Figure 4: (a-c) Constant-energy maps of the INS spectra in the plane for pure (left) and Li-MnTe (right) between 33 meV and 35 meV. The BZ is outlined by a thin grey line. The observed splitting in pure MnTe is indicated by red arrows. (d) Energy dependence of the local dynamic susceptibility extracted from INS data, with effective moments obtained from the total moment sum rule. (e) ARPES 100 meV integrated constant energy contour of Li-MnTe centered around the Fermi level, showing the first and second electronic BZ. (f) Second energy derivative high-symmetry cut along K--K with altermagnetic splitting of meV indicated by red bars.
@@
To understand whether the observed dispersion and lineshape for Li-MnTe can be explained by a persistent splitting which is simply broadened by the addition of itinerant carriers, we fit the data to two distinct LSWT models with refined , , : (1) with and zero, where splitting is necessarily absent; (2) with non-zero difference fixed at the value refined for MnTe. and were allowed to refine under this constraint [Fig. 3(f,g)].
The former scenario, which is identical to the “Li-doped MnTe center” presented earlier, underestimates both the spin wave energy and the linewidth near the expected splitting (Table S1).
@@
For the model with , we find improved agreement with experimental energies, further suggesting that the splitting in MnTe may survive Li-doping.